# VideoSAVi: Self-Aligned Video Language Models without Human Supervision

### v2

@@ 4.2 Ablation Studies
Table 6 illustrates the comparative efficiency of preference learning vs. parameter scaling for enhancing video understanding. The results show that VideoSAVi consistently delivers substantial improvements across all model sizes and benchmarks. Notably, a 1B parameter model with VideoSAVi (48.3 on LongVideoBench) outperforms a baseline 2B model (48.0), while a 2B model with VideoSAVi (52.3) surpasses a baseline 4B model (51.9). These findings confirm that addressing alignment issues through our self-supervised preference optimization yields greater performance gains than merely increasing model capacity. This efficiency advantage becomes particularly significant when considering the computational resources required for scaling parameters vs. our approach, which requires only self-generated preference pairs.
@@ 5 Conclusion
VideoSAVi is a novel self-aligned framework that employs a self-critiquing mechanism to detect and correct spatial and temporal reasoning failures in video-language models. It generates high-quality preference pairs directly from video content without requiring external supervision. Extensive experiments show substantial performance improvements over baseline models on a comprehensive set of benchmarks, demonstrating robust generalization across diverse test conditions. The method is computationally efficient and model-agnostic and shows remarkable parameter efficiency by enabling smaller models to outperform larger baselines. Future work will explore actor-critic frameworks for continuous self-improvement, further enhancing video understanding without reliance on external annotations or proprietary models.
@@
To ensure the reproducibility of our research, we rely on publicly available datasets and frameworks. Our training videos are sourced from open datasets, including Star (Wu et al., 2021), Vidal (Zhu et al., 2024), and WebVid (Bain et al., 2021). We implement our approach using the SWIFT training framework (Zhao et al., 2024) and evaluate it using the LMMs-Eval (Zhang et al., 2024a), both of which are open-source. Further, we will release our final trained model checkpoints and the complete set of model-generated preference pairs. A list of all prompts used in our implementation can be found in Section A.9.
@@ A.1 Iteration-wise Analysis
Figure 3 demonstrates the iteration-wise performance improvements of VideoSAVi across six video understanding benchmarks. Each iteration represents a complete cycle of our self-alignment pipeline, where the model (1) generates questions targeting specific reasoning capabilities and produces initial responses, (2) self-critiques these responses to identify reasoning errors, (3) creates revised responses based on critique feedback, and (4) uses the revised and initial responses to form preference pairs for DPO training. The consistent upward trajectory across all benchmarks validates our hypothesis that self-critiquing enables effective self-training without external supervision. Most notably, we observe substantial gains on tasks requiring complex reasoning: NeXTQA (+3.6%), MVBench (+5.1%), and EgoSchema (+6.8%). These improvements are particularly significant since EgoSchema features ego-centric videos with complex temporal relationships, while MVBench demands both fine-grained spatial understanding and temporal reasoning. The performance gains exhibit different patterns across benchmarks. TempCompass shows more modest improvements (+0.8%), likely because our baseline model already performs competitively on this benchmark. In contrast, PerceptionTest demonstrates steady improvement (+3.9%), indicating that VideoSAVi’s self-critiquing effectively addresses perceptual reasoning errors. For long-form videos (LongVideoBench), we observe a +2.0% improvement, confirming that our approach can enhance extended temporal reasoning capabilities. The iteration-wise analysis reveals that the most significant improvements occur in earlier iterations (especially iterations 1-2). This pattern suggests that VideoSAVi quickly identifies and corrects the most critical reasoning errors, followed by more minor refinements in subsequent iterations. Importantly, unlike methods that rely on external supervision or proprietary models, VideoSAVi achieves these improvements entirely through self-refinement, demonstrating the usability of self-aligned video understanding.
@@ A.2 Preference Composition Analysis
The decomposition of preference types in Table 7 reveals distinct task-specific effectiveness patterns. Spatial-only preferences lead to significant improvements on perception-intensive benchmarks (PerceptionTest: +3.3%, MVBench: +3.3%), demonstrating that targeting visual reasoning errors substantially enhances scene understanding capabilities. These gains align with the inherent spatial reasoning demands of these benchmarks, which require precise object localization and attribute recognition.
Temporal-only preferences demonstrate complementary strengths, excelling on benchmarks with extended temporal dependencies (EgoSchema: +4.1%, LongVideoBench: +1.4%, NeXTQA: +2.5%). The effectiveness on EgoSchema is particularly noteworthy, as egocentric videos contain complex action sequences requiring precise temporal ordering and causality inference. Temporal preference learning effectively addresses common temporal reasoning failures, including sequence ordering errors and causal misattributions.
The integration of both preference types yields strong results on several benchmarks. For MVBench, the full model achieves a +5.1% improvement, exceeding the sum of individual components (+3.3% spatial, +1.6% temporal). This non-linear gain suggests that joint optimization addresses compound reasoning errors that span both dimensions, particularly for tasks requiring spatiotemporal reasoning (e.g., tracking object state changes over time).
@@ A.3 Training Dataset Composition Analysis
To investigate the impact of dataset diversity on self-alignment performance, we conducted an ablation study using individual datasets while maintaining consistent pipeline parameters across four iterations. As shown in Table 8, VideoSAVi demonstrates dataset-agnostic improvements, achieving gains over the baseline InternVL2.5 model regardless of the source dataset. Star (Wu et al., 2021) videos particularly enhance complex reasoning benchmarks (NeXTQA +2.1%, EgoSchema +3.5%), likely due to their rich situated contexts. Vidal (Zhu et al., 2024) contributes balanced improvements across benchmarks, while WebVid (Bain et al., 2021) exhibits stronger contributions to perception-oriented tasks (PerceptionTest +2.8%). Notably, the performance gap between individual datasets and the combined approach (gap of 1.5-3.3% across benchmarks) demonstrates that VideoSAVi does not merely overfit to specific dataset preferences but rather benefits from diverse visual contexts. This finding reveals a critical insight: effective preference learning for video understanding requires exposure to varied temporal dynamics, visual compositions, and reasoning scenarios that single-source datasets cannot provide in isolation. The complementary nature of different video sources enables the model to develop more robust reasoning capabilities applicable across diverse benchmarks, confirming that dataset diversity serves as an implicit regularization factor in self-supervised preference optimization.
@@ A.5.1 Dataset Composition
    Training Distribution: 50 videos each from Star (Wu et al., 2021), WebVid (Bain et al., 2021), and Vidal (Zhu et al., 2024) datasets (150 total), representing the model’s original training domain.
@@
    Cross-Video: 50 additional unseen videos from the same datasets.
@@
    Full OOD: 40 videos from ActivityNet (Yu et al., 2019), representing a completely different domain with more complex activities.
For each video, we generate a rich set of questions using GPT-4o (Hurst et al., 2024) according to the prompt shown in Figure 13. This results in 8 questions per video (2 per category), totaling 1,920 question-answer pairs across all test conditions.
@@ A.5.2 Test Condition Categories
We design five distinct test conditions with progressively increasing difficulty:
@@
    Training Distribution. Standard format questions following typical video QA patterns applied to videos from the training distribution. These serve as our baseline for measuring generalization gaps.
@@
    Cross-Question. Questions that probe the same knowledge as Training Distribution but use difficult language, more complex syntax, and uncommon vocabulary. Examples include reformulating “What happens after X?” as “Describe the sequential events following the occurrence of X.” This tests generalization while keeping visual content consistent.
@@
    Cross-Video. Standard question formats applied to unseen videos from the same datasets. This tests visual generalization while keeping question patterns consistent.
@@
    Full OOD. Standard question formats applied to videos from entirely different domains (ActivityNet), testing both visual and domain generalization.
@@
    Adversarial. Questions deliberately designed to challenge video understanding systems by:
    • Directing attention to subtle background elements when foreground action is prominent.
    • Requiring reasoning about elements during moments of distraction.
    • Asking about negations or counterfactuals (e.g., “What does NOT change?”).
    • Focusing on elements briefly visible or partially occluded.
@@
    Compositional. Questions requiring integration of multiple reasoning types, such as:
    • Comparing temporal sequences across different spatial regions.
    • Tracking object state changes across discontinuous temporal segments.
    • Reasoning about causal chains involving multiple objects or actors.
    • Analyzing spatial transformations over time.
@@ A.9 Prompt Design for Self-Aligning Video-LLMs
Test Condition Question Generation Prompt for Generalization Assessment You will generate diverse question types to test a video understanding model’s generalization capabilities across increasingly challenging conditions. For each video, create questions in the following categories: Training Distribution Questions: Create standard format questions that follow typical patterns: “What object appears after the person sits down?” “What is the spatial relationship between [object A] and [object B]?” “How many people are visible in the scene?” Cross-Question Variants: Reformulate standard questions using novel phrasing and more complex language: “Describe the causal relationship between the sitting action and subsequent object appearance” “Elaborate on the configuration of entities in relation to the central figure” “What sequential patterns of movement can be discerned among the visible actors?” Adversarial Questions: Create questions with deliberately misleading cues or that require focusing on non-obvious elements: “What is the main action occurring in the background?” (when foreground action is prominent) “What subtle change occurs to the leftmost object while attention would naturally focus on the center?” “Ignoring the primary movement, what secondary action follows the initial event?” Compositional Questions: Formulate questions requiring multiple reasoning steps or integration of different reasoning types: “Compare the temporal order of actions on the left versus right sides of the frame” “How does the spatial configuration change from the beginning to the end of the sequence?” “What causal chain connects the initial object arrangement to the final state?” For each video, generate 8 questions (2 per category) with 4 answer choices and 1 correct choice. Ensure questions are answerable from the video content, specific, and aligned with their respective category definitions. For adversarial questions, identify what makes them challenging (e.g., distraction, subtlety, misdirection).

### v3

@@ 4.2 Ablation Studies
Table 5 illustrates the comparative efficiency of preference learning vs. parameter scaling for enhancing video understanding. The results show that VideoSAVi consistently delivers substantial improvements across all model sizes and benchmarks. While larger models generally achieve better performance, VideoSAVi offers complementary and cost-effective improvements. For instance, a 1B parameter model with VideoSAVi (48.3% on LongVideoBench) achieves comparable performance to a baseline 2B model (48.0%). This efficiency advantage becomes particularly significant when considering the computational resources required for scaling parameters vs. our approach, which requires only self-generated preference pairs.
Human Evaluation.
To validate the quality of self-generated preference pairs, we conduct a rigorous human evaluation with 6 external evaluators across 516 videos (261 temporal and 255 spatial). Evaluators are shown videos along model-generated questions and preference pairs without revealing the reasoning category. A pair is considered correct if and only if: (1) the question is answerable from video content, (2) the preferred response is accurate, and (3) the dispreferred response contains clear errors. This strict protocol reveals strong performance for both spatial relationships (71.3%) and temporal ordering (67.1%), with an overall quality of 69.2%. These results are particularly notable given that the preference learning process is fully self-supervised, operating without human intervention or external models. This confirms that our self-critiquing mechanism identifies genuine reasoning errors rather than arbitrary preferences, demonstrating that self-alignment can generate high-quality training signals for complex video understanding tasks.
Model-Agnostic Improvements.
Model TempCompass PerceptionTest NeXTQA MVBench EgoSchema LongVideoBench VideoLLaVA (Lin et al., 2024) 34.3 42.6 58.4 34.1 18.8 39.1 + VideoSAVi 36.8 +2.5 45.9 +3.3 63.2 +4.8 38.7 +4.6 23.5 +4.7 41.8 +2.7 LLaVA-NeXT-Interleave (Li et al., 2024b) 53.2 51.0 67.3 46.5 51.0 44.8 + VideoSAVi 55.6 +2.4 54.4 +3.4 71.2 +3.9 50.3 +3.8 54.6 +3.6 47.9 +3.1 LLaVA-OneVision (Li et al., 2024a) 64.1 57.5 79.3 56.7 64.0 56.3 + VideoSAVi 66.3 +2.2 60.8 +3.3 80.9 +1.6 59.9 +3.2 65.7 +1.7 58.7 +2.4 Qwen2-VL (Wang et al., 2024b) 68.9 62.1 75.6 64.9 59.2 55.6 +VideoSAVi 68.4 -0.5 65.4 +3.3 78.8 +3.2 68.3 +3.4 62.8 +3.6 58.1 +2.5 Qwen2.5VL (Bai et al., 2025) 72.3 68.6 75.8 65.2 60.9 60.7 +VideoSAVi 72.4 +0.1 68.9 +0.3 75.9 +0.1 65.4 +0.2 61.0 +0.1 60.7 +0.0 Qwen2.5-Omni (Xu et al., 2025) 69.6 66.6 75.2 63.0 61.8 56.9 +VideoSAVi 69.9 +0.3 66.9 +0.3 75.6 +0.4 64.3 +1.3 61.9 +0.1 57.2 +0.3
Table 7: Model-agnostic improvements from VideoSAVi.
Table 7 demonstrates the broad applicability of our approach across diverse architectures. VideoSAVi consistently enhances performance, delivering substantial improvements on nearly all benchmarks, regardless of the base model. The gains are particularly notable for VideoLLaVA (Lin et al., 2024), with improvements of +4.8 percentage points on NeXTQA, +4.7 on EgoSchema, and +4.6 on MVBench, showing that models with greater headroom benefit most from our preference optimization. For mid-tier models like LLaVA-NeXT-Interleave (Li et al., 2024b), VideoSAVi provides well-balanced improvements (+3.9 on NeXTQA, +3.8 on MVBench, +3.6 on EgoSchema), strengthening both spatial and temporal reasoning capabilities. Notably, VideoSAVi also enhances top-performing models, improving LLaVA-OneVision (Li et al., 2024a) by +1.6 on NeXTQA, reaching an impressive 80.9% and boosting Qwen2-VL (Wang et al., 2024b) across five benchmarks. For state-of-the-art models like Qwen2.5VL (Bai et al., 2025) and Qwen2.5-Omni (Xu et al., 2025) that have already undergone post-training refinements using DPO, VideoSAVi still delivers consistent but more modest improvements, with notable gains on MVBench (+0.2 and +1.3 respectively), demonstrating our method’s effectiveness even when applied to already-optimized models. These consistent cross-architecture improvements, averaging +2.1 percentage points across all models and benchmarks, confirm that our self-aligning preference learning framework addresses fundamental challenges in video understanding rather than exploiting model-specific characteristics.
Impact of Preference Data Size.
(a) TempCompass
(b) PerceptionTest
(c) NeXTQA
(d) MVBench
(e) EgoSchema
(f) LongVideoBench
Figure 2: Impact of preference data size on model performance.
Figure 2 demonstrates how VideoSAVi’s self-alignment pipeline consistently improves performance across all benchmarks as the percentage of preference pairs from the final (fourth) iteration of self-training increases. In contrast, TPO (Li et al., 2025b) shows minimal improvements or plateaus (particularly on EgoSchema), and Hound-DPO (Zhang et al., 2024b) experiences notable performance drops after initial improvements. VideoSAVi’s robustness and steady progress throughout training stems from its self-critiquing mechanism that targets challenging reasoning cases by effectively identifying subtle temporal inconsistencies and spatial reasoning errors. The steep improvement curve on EgoSchema (+6.8 percentage points) highlights how our approach excels at complex ego-centric understanding tasks, while the consistent gains on LongVideoBench demonstrate effective temporal alignment even when competing methods struggle with longer-form content.
Preservation of Core Capabilities.
To address concerns about potential overfitting during self-training, we evaluate VideoSAVi’s performance across a diverse set of benchmarks, assessing the backbone model’s pre-training performance (Chen et al., 2024). As shown in Table 8, VideoSAVi largely maintains or improves performance across all domains, preserving multimodal understanding (MMMU (Yue et al., 2024): +0.1 percentage points), document comprehension (DocVQA (Mathew et al., 2021): +0.6), and OCR reasoning (TextVQA (Mathew et al., 2021): +0.6), while enhancing video captioning (VideoChatGPT (Maaz et al., 2024): +0.27) and reducing hallucination (POPE (Li et al., 2023b): +0.2). This preservation is achieved by freezing the vision encoder and projector while applying low-rank LoRA adapters (rank 8) to update only 0.1% of language model parameters, retaining original visual feature extraction while enhancing language model reasoning over visual features.
Ability Benchmark Base +VideoSAVi
[…]
Reasoning GSM8k 77.4 76.9-$_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{-0.5}}}$
[…]
Table 8: VideoSAVi preserve core capabilities.
Model MVB ES LVB
InternVL2.5 69.8 52.0 57.8
+ Noisy Pref. 71.2+1.4 57.1+5.1 58.2+0.4
[…]
Table 9: Robustness to preference data noise. MVB: MVBench, ES: EgoSchema, LVB: LongVideoBench.
Robustness to Noisy Preferences.
To test VideoSAVi’s robustness to noisy preference data, we deliberately flip the labels of 30% of our 24,000 preference pairs, introducing substantial incorrect training signals. Table 9 shows that despite this significant noise, VideoSAVi with noisy preferences still consistently outperforms the baseline (+1.4 percentage points on MVBench, +5.1 on EgoSchema, +0.4 on LongVideoBench), while the full VideoSAVi pipeline maintains superior performance. This robustness stems from DPO’s inherent regularization through KL-divergence constraints and the on-policy nature of our self-generated preference pairs, which stay within the model’s natural output distribution and avoid instability associated with external preference data.
@@ 5 Conclusion
VideoSAVi is a novel self-aligned framework that employs a self-critiquing mechanism to detect and correct spatial and temporal reasoning failures in video-language models. It generates high-quality preference pairs directly from video content without requiring external supervision. Extensive experiments show substantial performance improvements over baseline models on a comprehensive set of benchmarks, demonstrating robust generalization across diverse test conditions. The method is computationally efficient and model-agnostic, enabling smaller models to outperform larger baselines through remarkable parameter efficiency. Future work will explore self-rewarding mechanisms and automatic hallucination detection frameworks, further enhancing video understanding reliability without reliance on external annotations or proprietary models.
@@
To ensure the reproducibility of our research, we rely on publicly available datasets and frameworks. Our training videos are sourced from open datasets, including Star (Wu et al., 2021), Vidal (Zhu et al., 2024), and WebVid (Bain et al., 2021). We implement our approach using the SWIFT training framework (Zhao et al., 2024) and evaluate it using the LMMs-Eval (Zhang et al., 2024a), both of which are open source. Further, we will release our final trained model checkpoints and the complete set of model-generated preference pairs. A list of all prompts used in our implementation can be found in Appendix §A.9.
@@ A.1 Iteration-wise Analysis
Figure 3 demonstrates the iteration-wise performance improvements of VideoSAVi across six video understanding benchmarks. Each iteration represents a complete cycle of our self-alignment pipeline, where the model (1) generates questions targeting specific reasoning capabilities and produces initial responses, (2) self-critiques these responses to identify reasoning errors, (3) creates revised responses based on critique feedback, and (4) uses the revised and initial responses to form preference pairs for DPO training. The consistent upward trajectory across all benchmarks validates our hypothesis that self-critiquing enables effective self-training without external supervision. Most notably, we observe substantial gains on tasks requiring complex reasoning: NeXTQA (+3.6 percentage points), MVBench (+5.1), and EgoSchema (+6.8). These improvements are particularly significant since EgoSchema features ego-centric videos with complex temporal relationships, while MVBench demands both fine-grained spatial understanding and temporal reasoning. The performance gains exhibit different patterns across benchmarks. TempCompass shows more modest improvements (+0.8), likely because our baseline model already performs competitively on this benchmark. In contrast, PerceptionTest demonstrates steady improvement (+3.9), indicating that VideoSAVi’s self-critiquing effectively addresses perceptual reasoning errors. For long-form videos (LongVideoBench), we observe a +2.0 improvement, confirming that our approach can enhance extended temporal reasoning capabilities. The iteration-wise analysis reveals that the most significant improvements occur in earlier iterations (especially iterations 1-2). This pattern suggests that VideoSAVi quickly identifies and corrects the most critical reasoning errors, followed by more minor refinements in subsequent iterations. Importantly, unlike methods that rely on external supervision or proprietary models, VideoSAVi achieves these improvements entirely through self-refinement, demonstrating the usability of self-aligned video understanding.
@@ A.2 Preference Composition Analysis
The decomposition of preference types in Table 10 reveals distinct task-specific effectiveness patterns. Spatial-only preferences lead to significant improvements on perception-intensive benchmarks (PerceptionTest: +3.3 percentage points, MVBench: +3.3), demonstrating that targeting visual reasoning errors substantially enhances scene understanding capabilities. These gains align with the inherent spatial reasoning demands of these benchmarks, which require precise object localization and attribute recognition.
Temporal-only preferences demonstrate complementary strengths, excelling on benchmarks with extended temporal dependencies (EgoSchema: +4.1, LongVideoBench: +1.4, NeXTQA: +2.5). The effectiveness on EgoSchema is particularly noteworthy, as egocentric videos contain complex action sequences requiring precise temporal ordering and causality inference. Temporal preference learning effectively addresses common temporal reasoning failures, including sequence ordering errors and causal misattributions.
The integration of both preference types yields strong results on several benchmarks. For MVBench, the full model achieves a +5.1 improvement, exceeding the sum of individual components (+3.3 spatial, +1.6 temporal). This non-linear gain suggests that joint optimization addresses compound reasoning errors that span both dimensions, particularly for tasks requiring spatiotemporal reasoning (e.g., tracking object state changes over time).
@@ A.3 Training Dataset Composition Analysis
To investigate the impact of dataset diversity on self-alignment performance, we conducted an ablation study using individual datasets while maintaining consistent pipeline parameters across four iterations. As shown in Table 11, VideoSAVi demonstrates dataset-agnostic improvements, achieving gains over the baseline InternVL2.5 model regardless of the source dataset. Star (Wu et al., 2021) videos particularly enhance complex reasoning benchmarks (NeXTQA +2.1 percentage points, EgoSchema +3.5), likely due to their rich situated contexts. Vidal (Zhu et al., 2024) contributes balanced improvements across benchmarks, while WebVid (Bain et al., 2021) exhibits stronger contributions to perception-oriented tasks (PerceptionTest +2.8). Notably, the performance gap between individual datasets and the combined approach (gap of 1.5-3.3 percentage points across benchmarks) demonstrates that VideoSAVi does not merely overfit to specific dataset preferences but rather benefits from diverse visual contexts. This finding reveals a critical insight: effective preference learning for video understanding requires exposure to varied temporal dynamics, visual compositions, and reasoning scenarios that single-source datasets cannot provide in isolation. The complementary nature of different video sources enables the model to develop more robust reasoning capabilities applicable across diverse benchmarks, confirming that dataset diversity serves as an implicit regularization factor in self-supervised preference optimization.
@@ A.5.1 Dataset Composition
    Training Distribution: 100 videos each from Star (Wu et al., 2021), WebVid (Bain et al., 2021), and Vidal (Zhu et al., 2024) datasets (300 total), representing the model’s original training domain.
@@
    Cross-Video: 150 additional unseen videos from the same datasets.
@@
    Full OOD: 150 videos from ActivityNet (Yu et al., 2019), representing a completely different domain with more complex activities.
For each video, we generate a rich set of questions using GPT-4o (Hurst et al., 2024) according to the prompt shown in Figure 13.
@@ A.5.2 Test Condition Categories
We design six distinct test conditions with progressively increasing difficulty:
@@
    Training Distribution. Standard format questions applied to training distribution videos. These serve as our baseline for measuring generalization gaps.
@@
    Cross-Question. Questions with complex syntax and uncommon vocabulary applied to training distribution videos. This tests linguistic generalization while keeping visual content consistent.
@@
    Cross-Video. Standard format questions applied to unseen videos from the same datasets. This tests visual generalization while keeping question patterns consistent.
@@
    Full OOD. Standard format questions applied to ActivityNet videos, testing both visual and domain generalization.
@@
    Adversarial. Deliberately challenging questions applied to training distribution videos, requiring focus on subtle or non-obvious elements.
@@
    Compositional. Multi-step reasoning questions applied to training distribution videos, requiring integration of multiple reasoning types.
@@ A.9 Prompt Design for Self-Aligning Video-LLMs
Test Condition Question Generation Prompt for Generalization Assessment You will generate diverse question types to test a video understanding model’s generalization capabilities across increasingly challenging conditions. For each video, create questions in the following categories: Training Distribution Questions: Create standard format questions that follow typical patterns: “What object appears after the person sits down?” “What is the spatial relationship between [object A] and [object B]?” “How many people are visible in the scene?” Cross-Question Variants: Reformulate standard questions using novel phrasing and more complex language: “Describe the causal relationship between the sitting action and subsequent object appearance” “Elaborate on the configuration of entities in relation to the central figure” “What sequential patterns of movement can be discerned among the visible actors?” Cross-Video Questions: Apply standard question formats to unseen videos from the same datasets: Use the same question templates as Training Distribution. Test visual generalization with consistent question complexity. Focus on transferring learned patterns to new visual content. Adversarial Questions: Create questions with deliberately misleading cues or that require focusing on non-obvious elements: “What is the main action occurring in the background?” (when foreground action is prominent) “What subtle change occurs to the leftmost object while attention would naturally focus on the center?” “Ignoring the primary movement, what secondary action follows the initial event?” Compositional Questions: Formulate questions requiring multiple reasoning steps or integration of different reasoning types: “Compare the temporal order of actions on the left versus right sides of the frame” “How does the spatial configuration change from the beginning to the end of the sequence?” “What causal chain connects the initial object arrangement to the final state?” For each video, generate 8 questions (2 per category) with 4 answer choices and 1 correct choice. Ensure questions are answerable from the video content, specific, and aligned with their respective category definitions. For adversarial questions, identify what makes them challenging (e.g., distraction, subtlety, misdirection).